OpenStudy (anonymous):

Using elimination, what are the values of x and y for the system of equations below? 2x+y=10 4x-y=8

OpenStudy (amistre64):

what are the steps to take in order to eliminate?

OpenStudy (anonymous):

im sorry i was hoping you would know. the possible answers are A.x=-9, y=28 B.x=-1, y=12 C.x=3,y=4 D.x=6,y=8

OpenStudy (amistre64):

i DO know ... but im not the one that needs to learn it :) i can hlep you out if you are stuck

OpenStudy (amistre64):

the elimination method is named that becasue we get rid of one of the variables ...

OpenStudy (amistre64):

and that can be done by adding the equations togther; but first we might need to modify one of them by multiplication

OpenStudy (amistre64):

if we add these equations togther; will one of the variable go away? 2x+y=10 4x-y=8

OpenStudy (anonymous):

the Y (i think)

OpenStudy (amistre64):

very good :) so lets add them and see .. 2x +y = 10 4x -y = 8 ---------- 6x +0 = 18 ; you wwere right :) now how does this make our life easier?

OpenStudy (anonymous):

do you subtract 18 from both sides or substitute something?

OpenStudy (amistre64):

not quite. its possible, but it doesnt get us anywhere. 6x = 18 ; this means that 6 times "something" equals 18 what does x have to equal ??

OpenStudy (anonymous):

3

OpenStudy (amistre64):

excellent; now we can stop there and check the answer opitions to narrow it down ...

OpenStudy (amistre64):

which answer has x=3 in it?

OpenStudy (anonymous):

haha its pretty obvious, C right

OpenStudy (amistre64):

yes :) but to dbl chk; lets see if y=4 will work out for us ... 2(3) +4 = 10 6 + 4 = 10 ... thats good 4(3) -4 = 8 12 -4 = 8 ; and its works for that one too ... id go with C :)

OpenStudy (anonymous):

Thank you! oh and im brand new to this site. How do i give you props? do i click the good answer thing?

OpenStudy (amistre64):

if you feel that you have been given a good answer, the good answer button is there for you to express it :) and a simple thank you goes along way too

OpenStudy (amistre64):

if you dont see a good answer button; you can refresh your browser by hitting the f5 key. :)

OpenStudy (anonymous):

ok good i did both

OpenStudy (anonymous):

thanks again

OpenStudy (amistre64):

youre quite welcome, good luck ;)